OpenStudy (anonymous):

Compete the pattern 243,81,27,9,3,--,--

OpenStudy (anonymous):

divide the next number by 3 (243/3=81. 81/3=27........)

OpenStudy (anonymous):

3/3=1, 1/3= one third, one third /3= one ninth